# Exploiting the diversity of modeling methods to probe systematic biases in strong lensing analyses

### v1

@@ 6.2 Recovery of source properties
Figure 4: Recovery of source properties among the six blind lens models shown in Fig. 2 (the legend in the right panel is also valid of other panels). Left panels: Azimuthally averaged two-point correlation functions $\xi(r)$ of the recovered sources, compared to the truth, which measures the correlation between pixels separated by a given distance $r$ (in arcsec). Middle panels: source effective radius $r_{\rm eff}$, computed within a circular aperture. Right panels: source axis ratio $q_{\rm s}$ estimated from the central moments of the source model. The bottom panel in each column indicates, the relative error with respect to to the true quantity, computed as $(\rm{truth}-\rm{model})/\rm{truth}$. For $r_{\rm eff}$ and $q_{\rm s}$, the gray shaded region indicates the mean and $1\sigma$ scatter among the models.
We show in Fig. 4 the recovery of several properties of the source galaxy: the two-point correlation function, the effective radius and the axis ratio. Bottom parts of each panel show the relative error computed as $(\rm{truth}-\rm{model})/\rm{truth}$ (i.e., negative values are over-estimates). We measure these properties within a square field of view of size $2\aas@@fstack{\prime\prime}2$, after projecting (using bi-cubic interpolation) each source model on a regular grid with 10 times higher resolution than the data. As shown in Fig. 2, such a field of view contains the entire flux of the source galaxy.
@@
We investigate the recovery of the size of the source galaxy through its effective radius $r_{\rm eff}$, that we define as the radius that encloses half of the total light within a circular aperture of radius $2\aas@@fstack{\prime\prime}2$. The middle panel of Fig. 4 shows $r_{\rm eff}$ and its relative error with respect to the true value for all source models. The effective radius is well recovered by all models with a maximum error of $2.5\%$. In addition we observe a tendency to over-estimate the effective radius, as only the Cluster+Exp model slightly under-estimates it, which is also the model with the lowest error. However, a better quantification of these errors should involve posterior distributions over the source models, which we do explore in this work. A first order uncertainty quantification can be obtained through the scatter among the different models, shown as the gray shaded area in Fig. 4. Given this scatter, the average source half-light radius remains lies close to $1\sigma$ from the true value.
@@ 6.3 Recovery of lens properties
To quantitatively compare the constraints on lens potential parameters among the six models, we show in Fig. 6 the joint posterior distributions for all mass model parameters, as well as true values from the data. Overall we find that the posterior distributions are within $\sim 3\sigma$ from the truth. The parameter with the smallest scatter relative to the posterior width is the logarithmic density slope $\gamma$. While all models are slightly biased towards values larger than the true value by approximately $2\%$, they are all compatible with it at $\lesssim 1\sigma$. Such a low scatter in the slope may be perhaps surprising, as it has been shown that the density slope may differ significantly between models (e.g., Etherington et al. 2022; Tan et al. 2023). However, we are in a regime where both the data and all models are parametrized by a single power-law density profile, hence the data is by construction a realization of the true model (modulo our inexact knowledge of the PSF and different numerics settings). Besides the density slope, the Einstein radius $\theta_{\rm E}$ also has low scatter among the models, which is expected as it is the primary quantity constrained by strong lensing observables. Models whose median values are further away from the truth also tend to have broader posteriors (larger uncertainties), which contribute to reduce the systematic bias (see also Fig. 7 below). The remaining mass model parameters all show visible scatter around the true values, while no systematic shift can be associated with a specific modeling method or model parameter.
@@
Figure 7: Uncertainty on a subset of key lens potential parameters, defined as the standard deviation of their marginalized posterior distribution (shown in Fig. 6): the Einstein radius $\theta_{\rm E}$, the mass density slope $\gamma$, the axis ratio $q$ and the external shear strength $\gamma_{\rm ext}$. The Einstein radius and the mass density slope are central quantities to many analyses, while the lens axis ratio and external shear are known to share degeneracies.
@@
From the full joint distributions over lens potential parameters, we investigate further the difference in uncertainties between the models. In Fig. 7 we plot posterior standard deviations for four parameters $\theta_{\rm E}$, $\gamma$, $q$ and $\gamma_{\rm ext}$. As expected from imaging lensing data, the uncertainty on the Einstein radius $\theta_{\rm E}$ is the smallest, with a relative precision on the order of $0.1\%$. For comparison, the relative precision on the mass density slope $\gamma$ is around $1.2\%$, and around to $1.5\%$ for the lens axis ratio $q$ (the relative error for $\gamma_{\rm ext}$ is inconclusive since it is close to zero). We notice that models with largest dynamic ranges (see Sect. 6.2) in their reconstructed source (Cluster+Exp+Lumweight) have smaller uncertainties on lens potential parameters. This trend is particularly clear for $\theta_{\rm E}$, $\gamma$, $q$ (first column in Fig. 7). Over the parameters shown in Fig. 7, the difference in uncertainties between the models remains relatively small, and amounts to a factor of approximately 1.6 between the least and most precise models.
Table 2: True values (first row) and summary statistics of source and lens properties inferred from the six lens models shown in Fig. 1. The second and third rows show the average, $\mu_{\rm model}$, and standard deviation, $\sigma_{\rm model}$, of the posterior means of all models. The fourth row gives the maximal bias among all models for each lens potential parameter. The two last rows correspond to the combined posterior distribution shown in black in Fig. 6.
Quantity $r_{\rm eff}$ [′′] $q_{\rm s}$ $m_{\rm s}$ [mag] $\theta_{\rm E}$ [′′] $\gamma$ $q_{\rm m}$ $\phi_{\rm m}$ $x_{\rm 0}$ [′′] $y_{\rm 0}$ [′′] $\gamma_{\rm ext}$ $\phi_{\rm ext}$ []
@@
Max. biasfootnotemark: $\times$ $\times$ $\times$ $1.7\sigma$ $1.6\sigma$ $-3.3\sigma$ $2.3\sigma$ $-5.5\sigma$ $-3.7\sigma$ $2.7\sigma$ $2.2\sigma$
@@
Biasfootnotemark: $\times$ $\times$ $\times$ $0.3\sigma$ $1.2\sigma$ $-0.6\sigma$ $-0.2\sigma$ $-1.6\sigma$ $-0.9\sigma$ $0.7\sigma$ $-0.3\sigma$
999 The source properties correspond to best-fit models, whereas lens properties relate to the corresponding marginalized posterior distributions shown in Fig. 6 (individual and combined models). Note that we do not provide combined estimates for the source properties as we do not have formal posterior distributions not uncertainty estimations for all source models.
@@
Ensemble models—namely, the combination of the posteriors of multiple models—can help improve modeling accuracy by correcting for the observed systematic biases of individual models. For this purpose, we show with black dotted-dashed lines and contours in Fig. 6 the result of combining the posteriors from each method, giving each model an equal weight. We find that these combined distributions are all within $1\sigma$ from the true values, except for the lens center along the $x$ direction which is at $1.7\sigma$. The marginalized statistics of the combined posterior are reported in the last row of Table 2 and compared to the simple average and standard deviation among individual models.
@@ 6.4 Correlations between the recovery of lens and source properties
Having in hand multiple lens models of the same data, we have the opportunity to explore how the accuracy of inferred lens and source properties correlate among the models. In particular, it is interesting to understand if certain biases observed in lens potential model parameters have an origin in biases in the reconstructed source light distribution, and vice versa. If such correlations exist, they could be used to design better parametrizations to jointly model the lens and source components, that specifically break these degeneracies. Such degeneracies may also be broken using non-lensing observations to further improve the accuracy of inferred lens and source properties (e.g., stellar kinematics of the source galaxy to place complementary constraints on its morphology). We emphasize that such correlations may be system- and data-dependent, which would warrant additional analyses complementary to ours.
@@ 7.1 Intrinsic source morphology
The spiral galaxy light profile that was used as the lensed source in producing the mock data presented in Sect. 4, has a prominent bright spot its center. The best-fit Sérsic index from the Sérsic+Shapelets model is approximately $1.6$, indicative of a cuspy radial profile. This feature, which we will refer to as the source cusp in the following paragraphs, consists of only a handful of pixels (¡20) that contain a significant amount of light ($\sim$5 %). This region is clearly hard to model, as shown by the reconstructions in the last two columns of Fig. 2, where only two models are able to capture it (Sérsic+Shapelets and Cluster+Exp+Lumweight). The remaining models fail to do so and leave behind characteristic residual flux at the data pixels where this compact region is multiply imaged.
@@
Here, we examine in more details how the performance of the model with the lowest brightness range (smoothest), Adaptive+Matérn, is affected by the prominence of the cusp. We select the central bright region with flux $>0.3$ and reduce the flux [^10] The noise map that is used as the covariance matrix in Eq. 3.1 is also changed accordingly. in each pixel by 50 and 95 %, creating two new mocks that we then model with exactly the same setup as the model shown in Fig. 2. The resulting (true) source and model residuals are shown in Fig. 8, while Fig. 9 shows the posterior distributions of the lens potential parameters. It can be seen that the more we suppress the cusp the better the model, i.e. we get less residual flux and less biased parameters.
@@ 7.2 Supersampling and inexact PSF model
Figure 10: Posterior distributions of lens potential parameters obtained with models exploring the role of the PSF. The green dot-dot-dashed line distributions correspond to the Cluster+Exp model (i.e., similar to the green model in Fig. 6) using the true supersampled PSF used for simulating the data (top left panel of Fig. 11). The red solid line distributions are showing, for reference, the blindly submitted model Cluster+Exp+Lumweight (i.e., the same model as in Fig. 6). The red dot-dashed and dotted distributions, also obtained with the Cluster+Exp+Lumweight model, use a supersampled (interpolated) version of the data-resolution PSF (bottom left panel in Fig. 11) or the true supersampled PSF, respectively. The main result of this comparison is that biases in lens model parameters are most reduced only with a combination of a more accurate PSF and a model that can capture magnified cuspy features in the source.
@@
We have seen in the previous section that a cuspy source can lead to biases in the lens potential parameters even in models that fit the central cusp well, in particular the Cluster+Exp+Lumweight model which achieved noise-level residuals. We now investigate whether using a supersampled PSF can further eliminate this bias. We implement PSF supersampling in two different ways: first, by interpolating in the pixel-level PSF (which is often implemented in practice when point sources are present in a lensed source); and second, by using the original supersampled PSF that was used to produce the mock data. Although the original PSF corresponds to a supersampling factor of 10 (i.e. each pixel is split into $10\times 10$ subpixels), this would be much too computationally expensive to employ directly. We therefore choose a supersampling factor of $f=5$ and downsample the original PSF with $f=10$ accordingly; this is what we will refer to as the “true supersampled PSF”. In the first model, we use bicubic interpolation in the pixel-level PSF to generate our interpolated $f=5$ supersampled PSF.
@@
It is interesting that without a luminosity-weighted source prior, the lens center coordinates are more accurately recovered (regardless of whether supersampling is used), despite all the other lens parameters being significantly biased. The Sérsic+Shapelet model produced a similar bias in the lens center coordinates as the luminosity-weighted models did, which is noteworthy since these are the only models that were able to reproduce the central cusp in the source well. While it is unclear exactly why this is the case, in real applications this may be ameliorated by the fact that the foreground lens light can furnish a prior in the lens center coordinates. Aside from this caveat, the bias in these parameters is at least somewhat reduced by supersampling with the true rather than interpolated PSF. We conclude that when fitting cuspy lensed sources, PSF supersampling can significantly reduce parameter biases only if it is accompanied by a regularization scheme (e.g. luminosity-weighted) that allows the source pixels to have steeper variations in the central bright region of the source galaxy where PSF supersampling is of the greatest benefit.
@@ 8.1 Quantifying model complexity and its impact on posterior uncertainties
In our work, as described in Sect. 3, one of the main differences between the modeling methods we consider is the way the source galaxy is reconstructed. In the past, several works focused on consistently comparing a set of lens models with different source reconstruction techniques developed under the same general formalism (i.e., the semi-linear inversion formalism, see Sect. 3.3, and Warren & Dye 2003), but differing in their regularization terms (e.g., Suyu et al. 2006; Tagore & Keeton 2014). Unfortunately, in the present work, there is no clear way to quantitatively and unambiguously rank the complexity of the source models considered here, given their fundamental differences in terms of mathematical formalism and underlying assumptions about the morphology of galaxies. One possibility would be to count their number of degrees of freedom, but this quantity is not readily accessible for all models. In regularized pixelated source models, the effective number of degrees of freedom is lower than them total number of source pixels, as regularization correlates source pixels over different spatial scales (e.g., Suyu et al. 2006; Nightingale & Dye 2015). As a concrete example, regularization based on sparsity in a wavelet basis do not allow to analytically and unambiguously measure the number of degrees of freedom as it imposes the sparsity simultaneously at various spatial scales with non-linear dependency on the lens deflection field.

### v2

@@ 6.2 Recovery of source properties
Figure 4: Recovery of source properties among the six blind lens models shown in Fig. 2 (the legend in the right panel is also valid of other panels). Left panels: Azimuthally averaged two-point correlation functions, $\xi(r)$, of the recovered sources, compared to the truth, which measures the correlation between pixels separated by a given distance, $r$ (in arcsec). Middle panels: Source effective radius $r_{\rm eff}$, computed within a circular aperture. Right panels: Source axis ratio, $q_{\rm s}$, estimated from the central moments of the source model. The bottom panel in each column indicates, the relative error with respect to to the true quantity, computed as $(\rm{truth}-\rm{model})/\rm{truth}$. For $r_{\rm eff}$ and $q_{\rm s}$, the shaded gray region indicates the mean and $1\sigma$ scatter among the models.
We show in Fig. 4 the recovery of several properties of the source galaxy: the two-point correlation function, the effective radius and the axis ratio. Bottom parts of each panel show the relative error computed as $(\rm{truth}-\rm{model})/\rm{truth}$ (i.e., negative values are over-estimates). We measure these properties within a square field of view of size $2\aas@@fstack{\prime\prime}2$, after projecting (using bi-cubic interpolation) each source model on a regular grid with 10 times higher resolution than the data. As is shown in Fig. 2, such a field of view contains the entire flux of the source galaxy.
@@
We investigate the recovery of the size of the source galaxy through its effective radius, $r_{\rm eff}$, which we define as the radius that encloses half of the total light within a circular aperture of radius $2\aas@@fstack{\prime\prime}2$. The middle panel of Fig. 4 shows $r_{\rm eff}$ and its relative error with respect to the true value for all source models. The effective radius is well recovered by all models with a maximum error of $2.5\%$. In addition we observe a tendency to over-estimate the effective radius, as only the Cluster+Exp model slightly under-estimates it, which is also the model with the lowest error. However, a better quantification of these errors should involve posterior distributions over the source models, which we do explore in this work. A first order uncertainty quantification can be obtained through the scatter among the different models, shown as the shaded gray area in Fig. 4. Given this scatter, the average source half-light radius remains lies close to $1\sigma$ from the true value.
@@ 6.3 Recovery of lens properties
To quantitatively compare the constraints on lens potential parameters among the six models, we show in Fig. 6 the joint posterior distributions for all mass model parameters, as well as true values from the data. Overall we find that the posterior distributions are within $\sim 3\sigma$ from the truth. The parameter with the smallest scatter relative to the posterior width is the logarithmic density slope $\gamma$. While all models are slightly biased towards values larger than the true value by approximately $2\%$, they are all compatible with it at $\lesssim 1\sigma$. Such a low scatter in the slope may be perhaps surprising, as it has been shown that the density slope may differ significantly between models (e.g., Etherington et al. 2022; Tan et al. 2024). However, we are in a regime where both the data and all models are parametrized by a single power-law density profile, hence the data is by construction a realization of the true model (modulo our inexact knowledge of the PSF and different numerics settings). Besides the density slope, the Einstein radius, $\theta_{\rm E}$, also has low scatter among the models, which is expected as it is the primary quantity constrained by strong lensing observables. Models whose median values are further away from the truth also tend to have broader posteriors (larger uncertainties), which contribute to reduce the systematic bias (see also Fig. 7 below). The remaining mass model parameters all show visible scatter around the true values, while no systematic shift can be associated with a specific modeling method or model parameter.
@@
Figure 7: Uncertainty on a subset of key lens potential parameters, defined as the standard deviation of their marginalized posterior distribution (shown in Fig. 6): the Einstein radius, $\theta_{\rm E}$, the mass density slope, $\gamma$, the axis ratio $q$, and the external shear strength $\gamma_{\rm ext}$. The Einstein radius and the mass density slope are central quantities to many analyses, while the lens axis ratio and external shear are known to share degeneracies.
@@
From the full joint distributions over lens potential parameters, we investigate further the difference in uncertainties between the models. In Fig. 7 we plot posterior standard deviations for four parameters, $\theta_{\rm E}$, $\gamma$, $q$, and $\gamma_{\rm ext}$. As is expected from imaging lensing data, the uncertainty on $\theta_{\rm E}$ is the smallest, with a relative precision on the order of $0.1\%$. For comparison, the relative precision on the mass density slope $\gamma$ is around $1.2\%$, and around to $1.5\%$ for the lens axis ratio $q$ (the relative error for $\gamma_{\rm ext}$ is inconclusive since it is close to zero). We notice that models with largest dynamic ranges (see Sect. 6.2) in their reconstructed source (Cluster+Exp+Lumweight) have smaller uncertainties on lens potential parameters. This trend is particularly clear for $\theta_{\rm E}$, $\gamma$, $q$ (first column in Fig. 7). Over the parameters shown in Fig. 7, the difference in uncertainties between the models remains relatively small, and amounts to a factor of approximately 1.6 between the least and most precise models.
Table 2: Posterior statistics of source and lens properties inferred from the six lens models shown in Fig. 1.
Quantity $r_{\rm eff}$ [′′] footnotemark: $q_{\rm s}$ footnotemark: $m_{\rm s}$ [mag] $\theta_{\rm E}$ [′′] $\gamma$ $q_{\rm m}$ $\phi_{\rm m}$ $x_{\rm 0}$ [′′] $y_{\rm 0}$ [′′] $\gamma_{\rm ext}$ $\phi_{\rm ext}$ []
@@
Max. bias††††footnotemark: †† $\times$ $\times$ $\times$ $1.7\sigma$ $1.6\sigma$ $-3.3\sigma$ $2.3\sigma$ $-5.5\sigma$ $-3.7\sigma$ $2.7\sigma$ $2.2\sigma$
@@
Bias††††footnotemark: †† $\times$ $\times$ $\times$ $0.3\sigma$ $1.2\sigma$ $-0.6\sigma$ $-0.2\sigma$ $-1.6\sigma$ $-0.9\sigma$ $0.7\sigma$ $-0.3\sigma$
999 In the second and third row $\mu_{\rm model}$ and $\sigma_{\rm model}$ show the average and standard deviation of the posterior means of all models, respectively. The fourth row gives the maximal bias among all models for each lens potential parameter. The two last rows correspond to the combined posterior distribution shown in black in Fig. 6.
@@
[…]
Ensemble models—namely, the combination of the posteriors of multiple models—can help improve modeling accuracy by correcting for the observed systematic biases of individual models. Depending on the assumptions about the original individual models, in particular regarding their statistical independence, there exist different approaches to combine their posteriors together. Here we follow a conservative approach and simply combine individual posteriors with equal weights. We show with dash-dotted black lines and contours in Fig. 6 the resulting combined posteriors. We find that these combined distributions are all within $1\sigma$ from the true values, except for the lens center along the $x$ direction which is at $1.7\sigma$. The marginalized statistics of the combined posterior are reported in the last row of Table 2 and compared to the simple average and standard deviation among individual models.
@@ 6.4 Correlations between the recovery of lens and source properties
Having in hand multiple lens models of the same data, we have the opportunity to explore how the accuracy of inferred lens and source properties correlate among the models. In particular, it is interesting to understand if certain biases observed in lens potential model parameters have an origin in biases in the reconstructed source light distribution, and vice versa. If such correlations exist, they could be used to design better parametrizations to jointly model the lens and source components that specifically break these degeneracies. Such degeneracies may also be broken using non-lensing observations to further improve the accuracy of inferred lens and source properties (e.g., stellar kinematics of the source galaxy to place complementary constraints on its morphology). We emphasize that such correlations may be system- and data-dependent, which would warrant additional analyses complementary to ours.
@@ 7.1 Intrinsic source morphology
The spiral galaxy light profile that was used as the lensed source in producing the mock data presented in Sect. 4, has a prominent bright spot its center. The best-fit Sérsic index from the Sérsic+Shapelets model is approximately $1.6$, indicative of a cuspy radial profile. This feature, which we shall refer to as the source cusp in the following paragraphs, consists of only a handful of pixels (¡20) that contain a significant amount of light ($\sim$5 %). This region is clearly hard to model, as is shown by the reconstructions in the last two columns of Fig. 2, where only two models are able to capture it (Sérsic+Shapelets and Cluster+Exp+Lumweight). The remaining models fail to do so and leave behind characteristic residual flux at the data pixels where this compact region is multiply imaged.
@@
Here, we examine in more details how the performance of the model with the lowest brightness range (smoothest), Adaptive+Matérn, is affected by the prominence of the cusp. We select the central bright region with flux $>0.3$ and reduce the flux [^10] The noise map that is used as the covariance matrix in Eq. 3.1 is also changed accordingly. in each pixel by 50 and 95 %, creating two new mocks that we then model with exactly the same setup as the model shown in Fig. 2. The resulting (true) source and model residuals are shown in Fig. 8, while Fig. 9 shows the posterior distributions of the lens potential parameters. It can be seen that the more we suppress the cusp the better the model; that is, we get less residual flux and less biased parameters.
@@ 7.2 Supersampling and inexact PSF model
Figure 10: Posterior distributions of lens potential parameters obtained with models exploring the role of the PSF. The dash-dotted green line distributions correspond to the Cluster+Exp model (i.e., similar to the green model in Fig. 6) using the true supersampled PSF used for simulating the data (top left panel of Fig. 11). The solid red line distributions are showing, for reference, the blindly submitted model Cluster+Exp+Lumweight (i.e., the same model as in Fig. 6). The dash-dotted and dotted red distributions, also obtained with the Cluster+Exp+Lumweight model, use a supersampled (interpolated) version of the data-resolution PSF (bottom left panel in Fig. 11) or the true supersampled PSF, respectively. The main result of this comparison is that biases in lens model parameters are most reduced only with a combination of a more accurate PSF and a model that can capture magnified cuspy features in the source.
@@
We have seen in the previous section that a cuspy source can lead to biases in the lens potential parameters even in models that fit the central cusp well, in particular the Cluster+Exp+Lumweight model which achieved noise-level residuals. We now investigate whether using a supersampled PSF can further eliminate this bias. We implement PSF supersampling in two different ways: first, by interpolating in the pixel-level PSF (which is often implemented in practice when point sources are present in a lensed source); and second, by using the original supersampled PSF that was used to produce the mock data. Although the original PSF corresponds to a supersampling factor of 10 (i.e., each pixel is split into $10\times 10$ subpixels), this would be much too computationally expensive to employ directly. We therefore choose a supersampling factor of $f=5$ and downsample the original PSF with $f=10$ accordingly; this is what we shall refer to as the “true supersampled PSF.” In the first model, we use bicubic interpolation in the pixel-level PSF to generate our interpolated $f=5$ supersampled PSF.
@@
It is interesting that without a luminosity-weighted source prior, the lens center coordinates are more accurately recovered (regardless of whether supersampling is used), despite all the other lens parameters being significantly biased. The Sérsic+Shapelet model produced a similar bias in the lens center coordinates as the luminosity-weighted models did, which is noteworthy since these are the only models that were able to reproduce the central cusp in the source well. While it is unclear exactly why this is the case, in real applications this may be ameliorated by the fact that the foreground lens light can furnish a prior in the lens center coordinates. Aside from this caveat, the bias in these parameters is at least somewhat reduced by supersampling with the true rather than interpolated PSF. We conclude that when fitting cuspy lensed sources, PSF supersampling can significantly reduce parameter biases only if it is accompanied by a regularization scheme (e.g., luminosity-weighted) that allows the source pixels to have steeper variations in the central bright region of the source galaxy where PSF supersampling is of the greatest benefit.
@@ 8.1 Quantifying model complexity and its impact on posterior uncertainties
In our work, as is described in Sect. 3, one of the main differences between the modeling methods we consider is the way the source galaxy is reconstructed. In the past, several works focused on consistently comparing a set of lens models with different source reconstruction techniques developed under the same general formalism (i.e., the semi-linear inversion formalism, see Sect. 3.3, and Warren & Dye 2003), but differing in their regularization terms (e.g., Suyu et al. 2006; Tagore & Keeton 2014). Unfortunately, in the present work, there is no clear way to quantitatively and unambiguously rank the complexity of the source models considered here, given their fundamental differences in terms of mathematical formalism and underlying assumptions about the morphology of galaxies. One possibility would be to count their number of degrees of freedom, but this quantity is not readily accessible for all models. In regularized pixelated source models, the effective number of degrees of freedom is lower than them total number of source pixels, as regularization correlates source pixels over different spatial scales (e.g., Suyu et al. 2006; Nightingale & Dye 2015). As a concrete example, the regularization based on sparsity and wavelets do not allow one to unambiguously estimate the number of degrees of freedom, because such regularization imposes sparsity simultaneously over various spatial scales.